. A chemieal engineer in studying the two renections shown in the table below In each ease, he fila a reaction venel with some mixture of the reactants and products at a constant tomperature of RN0 C and constant total pressure. Then, he measures t reaction entholpy All and reaction entropy As8 of the first reaction enthalpy AH and reaction free energy AG of the second reaction. The results of his measurements are ahown in the table he Complete the table. That is, caleulate AGG for the first reaction and AS for the eecond (Hound your answer to vero decimal places) Then, decide whether, under the conditions the engineer has set up, the reaction is spontaneous, the reverse reaction is apoianeous, ot neither forwand nor reverse reaction is spontaneous because the systen in at eqpuilibrium AH 951. kJ 2577. kJ Which is spontaneous? this reaction O the reverse reaction 0 neither as- Which is spontaneous? 0 this reaction O the reverse reaction O neither ooo: 25
Answer #1

1)

Al2O3(s) + 3H2(g) ------------------- 2 Al(s) + 3 H2O(g)

\DeltaH = 951 KJ                      T= 88.0C = 88.0+273=361 K

\DeltaS = 2577 J/K

\DeltaG = \Delta H - T\DeltaS

\DeltaG = 951 - 361x2577 = 20703 J

\DeltaG = 20.703 KJ

\DeltaG is positve sign for forword reaction.In the forword direction it is non-spontaneous

Hence , the reverse reaction is spontaneous.

2)

2 H2(g) + O2(g) ----------------- 2H2O(g)

\DeltaH = - 484 KJ

\DeltaG = - 25 KJ

T= 361K

\DeltaG = \Delta H - T \Delta S

\DeltaS = \Delta H - \Delta G/T = -484 - ( 25)/361 = - 1271.47 J/k

\DeltaS = - 1271.47 J/K

\DeltaG is negative sign. so the given equation is spontaneous.